# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ D-A2 $1>\frac{m}{\theta NT\log_{2}e}$
$=$ $\left(\frac{1}{2}\right)^{m-1}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\gamma\Omega}}\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right),$ (116b)
where Eq. (116a) holds because $\frac{x}{\frac{1}{\gamma}+x}$ an increasing function for $x\in\big{[}\frac{1}{\gamma},\infty\big{)}$.
By substituting Eq. (116b) into Eq. (113), we obtain
@@
$=\frac{m}{\theta NT\log_{2}e},$ (117b)
where Eq. (117a) holds because according to [49], for $s<0$,
$\lim_{x\to 0}\frac{\Gamma(s,x)}{x^{s}}=-\frac{1}{s}.$ (118)
D-A3 $1=\frac{m}{\theta NT\log_{2}e}$
For $1=\frac{m}{\theta NT\log_{2}e}$, $\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)=E_{1}\left(\frac{2m}{\Omega\gamma}\right)$. By substituting $\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)=E_{1}\left(\frac{2m}{\Omega\gamma}\right)$ and Eq. (116b) into Eq. (113), we obtain
@@ D-A3 $1=\frac{m}{\theta NT\log_{2}e}$
Eq. (119a) holds because according to [50]
$E_{1}(x)>\frac{1}{2}e^{-x}\ln\left(1+\frac{2}{x}\right),\quad x>0.\vspace{-2mm}$ (120)
Now we have completed the proof for the upper bound of $s_{\infty}$ for $m\geq 1$.
D-B Lower bound of $s_{\infty}$ for $m\geq 1$
Since $\sqrt{1-x}\leq 1$ with $x\in(0,1)$, the lower bound of $s_{\infty}$ in the Nakagami-$m$ fading channel is given by
@@ D-B Lower bound of $s_{\infty}$ for $m\geq 1$
Eq. (121b) holds since $m\geq 1$ and $\frac{1}{\gamma}+x>x$ for $x>0$. Since $\frac{m}{\Omega}$ is a finite constant, $\lim\limits_{\gamma\to+\infty}\frac{m}{\Omega\gamma}=0$.
D-B1 $1<\frac{m}{\theta NT\log_{2}e}$
$\Gamma(s,x)<\frac{(x+v_{s})^{s}-x^{s}}{v_{s}s}e^{-x},$ (122)
where $v_{s}=\Gamma(s+1)^{\frac{1}{s-1}}>0$. For simplicity, we let $c_{7}=\Gamma(m+1-\theta NT\log_{2}e)^{\frac{1}{m-1-\theta NT\log_{2}e}}$. Based on Eq. (122), for $m-\theta NT\log_{2}e\in(0,1]\cup[2,\infty)$, we obtain
@@ D-B1 $1<\frac{m}{\theta NT\log_{2}e}$
Eq. (123a) holds because $\lim\limits_{\gamma\to+\infty}\frac{\ln\left[\left(1+\frac{c_{7}\Omega}{m}\gamma\right)^{m-\theta NT\log_{2}e}-1\right]}{(m-\theta NT\log_{2}e)\ln\left(1+\frac{c_{7}\Omega}{m}\gamma\right)}=1$ with $c_{7}>0$ and $m-\theta NT\log_{2}e>0$. Eq. (123b) holds because $\lim\limits_{\gamma\to+\infty}\frac{\ln\left(1+\frac{c_{7}\Omega}{m}\gamma\right)}{\ln\gamma}=1$.
For $m-\theta NT\log_{2}e\in(1,2)$, we focus on the integral in Eq. (121a). When $m-\theta NT\log_{2}e\in(1,2)$, $\theta NT\log_{2}e<m$ holds. Thus, we find that $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ is upper bounded by a finite constant according to Eq. (114). Thus, we obtain that
$s_{\infty}\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}-\lim_{\gamma\to+\infty}\frac{\beta}{N\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}=1.$ (124)
D-B2 $1\geq\frac{m}{\theta NT\log_{2}e}$
For $m-\theta NT\log_{2}e<0$, we have
@@ D-B2 $1\geq\frac{m}{\theta NT\log_{2}e}$
$=\frac{m}{\theta NT\log_{2}e},$
where Eq. (125a) holds according to Eq. (118).
D-B3 $1=\frac{m}{\theta NT\log_{2}e}$
For $m-\theta NT\log_{2}e=0$, $\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right)=E_{1}\left(\frac{m}{\Omega\gamma}\right)$. According to Eq. (121), we have
@@ D-B3 $1=\frac{m}{\theta NT\log_{2}e}$
Eq. (126a) holds because according to [50]
@@
Therefore, according to the squeeze theorem, we have completed the proof for $m\geq 1$ based on the results in Eqs. (115), (117), (119), (123), (125), and (126).
@@
For $10.5\leq m<$ 1, there are some differences in the proof, which mainly lie in the proof of the upper bound with $1>\frac{m}{\theta NT\log_{2}e}$ and the lower bound.
For the upper bound with $1\leq\frac{m}{\theta NT\log_{2}e}$, the proof is same with Eqs. (115) and (119). For the upper bound with $1>\frac{m}{\theta NT\log_{2}e}$, Eq. (116) is changed into
@@
Then, by substituting Eq. (128) into Eq. (113), we obtain
$=\frac{m}{\theta NT\log_{2}e}.$
Eq. (129a) holds because $m-\theta NT\log_{2}e<0$ so that Eq. (118) holds.
D-D Lower Bound of $s_{\infty}$ for $m\in[0.5,1)$
Then, we focus on the lower bound of $s_{\infty}$ with $10.5\leq m<$ 1. The essence of this part of the proof is still the scaling of $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$. According to Eq. (121), the lower bound of $s_{\infty}$ can be expressed as
D-D1 $1<\frac{m}{\theta NT\log_{2}e}$
For $1<\frac{m}{\theta NT\log_{2}e}$, we transform Eq. (130) into
@@ D-D1 $1<\frac{m}{\theta NT\log_{2}e}$
Eq. (131) holds because $-\theta NT\log_{2}e<0$ and $\frac{1}{\gamma}+x>x$. As we mentioned before, $\int_{0}^{\infty}z^{a}e^{-x}{\rm d}x<\infty$ holds when $a>-1$. Since in this case $1<\frac{m}{\theta NT\log_{2}e}$, $-\theta NT\log_{2}e+m-1>-1$. Thus, in this case $\int_{0}^{\infty}x^{-\theta NT\log_{2}e+m-1}e^{-\frac{m}{\Omega}x}{\rm d}x=\left(\frac{\Omega}{m}\right)^{m-\theta NT\log_{2}e}\int_{0}^{\infty}t^{m-\theta NT\log_{2}e-1}e^{-t}{\rm d}t$ is a finite constant. Then, we obtain
$s_{\infty}\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}-\frac{1}{N}\lim_{\gamma\to+\infty}\frac{\beta}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}=1.$ (132)
D-D2 $1\geq\frac{m}{\theta NT\log_{2}e}$
For $1\geq\frac{m}{\theta NT\log_{2}e}$, we find
@@ D-D2 $1\geq\frac{m}{\theta NT\log_{2}e}$
$=\gamma^{\theta NT\log_{2}e-m}e^{\frac{m}{2\Omega\gamma}}K_{0}\left(\frac{m}{2\Omega\gamma}\right),$
where $K_{v}(z)$ is the modified Bessel function of the second kind [50]. Eq. (133a) holds because in this case $m-\frac{1}{2}\geq 0$ and $\frac{x}{\frac{1}{\gamma}+x}\leq 1$. According to [50], for the modified Bessel function of the second kind $K_{v}(z)$ with $v=0$, we have
$\lim_{z\to 0}\frac{K_{0}(z)}{-\ln(z)}=c_{8},$ (134)
where $c_{8}>0$ is a constant. Based on this conclusion, we transformed Eq. (133) into
@@
$=\frac{m}{\theta NT\log_{2}e},$ (136b)
where Eq. (136b) holds according to Eq. (134). According to the squeeze theorem, we have completed the proof for $m\in[0.5,1)$. Now we have completed the proof.
@@ Appendix E Proof of Theorem 6
With the $\Psi(\gamma)$ satisfies Theorem 3, Eq. (111) still holds. For the AWGN channel, we can easily obtain
@@
For an FBL-SISO system in a Nakagami-$m$ fading channel, we will prove that the following two inequalities hold.
In Eq. (139), $c_{3}>0$ is an arbitrary constant. Since $c_{3}>0$ can be an arbitrarily small constant, we obtain Eq. (49) based on Eqs. (138) and (139).
The derivation of Eq. (138) is based on the derivation of the lower bounds of $s_{\infty}$ in the proof of Theorem 5. For $m\geq 1$, we can directly obtain Eq. (138) according to Eqs. (123a), (125a), and (126b). For $m\in[0.5,1)$, we can obtain Eq. (138) according to Eqs. (132) and (136a). Thus, we omit the details here.
The derivation of Eq. (139) is based on the derivation of the upper bounds of $s_{\infty}$. Different from the proof of Theorem 5, we start the proof with the help of $\sqrt{1-x}\geq 1-\sqrt{x}$ with $x\in(0,1)$. We have
Based on $\frac{x}{x+1}\leq\ln(1+x)$ for $x>0$, we have $\frac{1}{1+\gamma x}\leq\ln\left(1+\frac{1}{\gamma x}\right)$. Thus, we transform Eq. (140) into
The integral in the upper bound of $s_{\infty}$ is given by
Note that, with $\gamma\to+\infty$, $\beta\to+\infty$. Thus, $01-m-\theta T\beta<$ 0. We obtain
$=$ $\left(\frac{c_{9}}{c_{9}+1}\right)^{\theta T\beta+m-1}\left(\frac{m}{\Omega}\right)^{\theta N\log_{2}e-m}e^{\frac{m}{\Omega\gamma}}\Gamma\left(m-\theta NT\log_{2}e,\frac{m(1+c_{9})}{\Omega\gamma}\right),$
where $c_{9}$ is an arbitrarily positive constant.
Based on this result, we obtain
For $1>\frac{m}{\theta NT\log_{2}e}$, by combining Eq. (118) with Eq. (144), we have
Note that, since $c_{9}$ is an arbitrarily positive constant, $\ln\left(1+\frac{1}{c_{9}}\right)$ is an arbitrarily positive constant.
The first part is $1\leq\frac{m}{\theta NT\log_{2}e}\leq 1+\frac{1}{\theta NT\log_{2}e}$ and $\frac{m}{\theta NT\log_{2}e}\geq 1+\frac{2}{\theta NT\log_{2}e}$, which means $0\leq m-\theta NT\log_{2}e\leq 1$ and $m-\theta NT\log_{2}e\geq 2$. We will utilize the conclusion that $\Gamma(s,x)\geq e^{-x}(1+x)^{s-1}$ for $x>0$ with $s\leq 1$ and $s\geq 2$. This can be easily proved by Jensen’s inequality. $(t+x)^{s-1}$ is a convex function of $t$ with a $s\leq 1$ and $s\geq 2$. Thus, we have
$\int_{x}^{\infty}y^{s-1}e^{-y}{\rm d}y=e^{-x}\int_{0}^{\infty}(t+x)^{s-1}e^{-t}{\rm d}t\geq e^{-x}\left(\int_{0}^{\infty}(t+x)e^{-t}{\rm d}t\right)^{s-1}=e^{-x}(1+x)^{s-1}.$ (146)
Now we have proved that $\Gamma(s,x)\geq e^{-x}(1+x)^{s-1}$ holds for $s\leq 1$ and $s\geq 2$.
Thus, by substituting Eq. (146) into Eq. (144), we have
@@
The first part is $1+\frac{1}{\theta NT\log_{2}e}<\frac{m}{\theta NT\log_{2}e}<1+\frac{2}{\theta NT\log_{2}e}$, which means $21<m-\theta NT\log_{2}e<$ 2. For $\Gamma(s,x)$ with $21<s<$ 2, according to [51], we have
$\Gamma(s,x)>\frac{(x+v_{s})^{s}-x^{s}}{v_{s}s}e^{-x},$ (148)
where $v_{s}$ is defined after Eq. (122). By substituting Eq. (148) into Eq. (144), we can obtain Eq. (149) with the similar procedures as shown in Eq. (123).
@@
Now we have completed the proof of Eq. (139).

### v4

@@ D-A Upper bound of $s_{\infty}$ for $m\geq 1$
$=$ $\left(\frac{1}{2}\right)^{m-1}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\gamma\Omega}}\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right).$ (118b)
@@
$=\frac{m}{\theta NT\log_{2}e}.$ (119b)
@@
$s_{\infty}$ $\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}-\lim_{\gamma\to+\infty}\frac{\beta}{N\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}$ (127)
$=1.$
@@
$=\frac{m}{\theta NT\log_{2}e}.$
D-A1 $1<\frac{m}{\theta NT\log_{2}e}$
For $1<\frac{m}{\theta NT\log_{2}e}$, we find that Eq. (116), which is presented at the bottom of this page, holds.
According to [14], $0<\int_{0}^{\infty}z^{a}e^{-x}{\rm d}x<\infty$ if $a>-1$. Since in this case $m>\theta NT\log_{2}e$, we obtain $m-\theta NT\log_{2}e-1>-1$. Therefore, the integral $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ is bounded by a finite constant. Thus, under the condition $1<\frac{m}{\theta NT\log_{2}e}$, Eq. (115) can be further expressed as
$s_{\infty}$ $\leq\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}$ (117)
$=1.$
Eq. (117) holds because
$\lim\limits_{\gamma\to+\infty}\frac{\log_{2}N}{N}=\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\Psi(\gamma)}{2\Psi(\gamma)}=0.$
is valid under the condition that $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$.
D-A2 $1>\frac{m}{\theta NT\log_{2}e}$
For $1>\frac{m}{\theta NT\log_{2}e}$, we focus on the integral
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x.$
This yields Eq. (118), which is shown at the bottom of this page. Eq. (118a) holds because $\frac{x}{\frac{1}{\gamma}+x}$ an increasing function for $x\in\big{[}\frac{1}{\gamma},\infty\big{)}$.
By substituting Eq. (118b) into Eq. (115), we arrive at Eq. (119), which is presented at the bottom of this page. Eq. (119a) holds because according to [47], the following inequality holds for $s<0$:
$\lim_{x\to 0}\frac{\Gamma(s,x)}{x^{s}}=-\frac{1}{s}.$ (120)
D-A3 $1=\frac{m}{\theta NT\log_{2}e}$
For $1=\frac{m}{\theta NT\log_{2}e}$, we have
$\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)=E_{1}\left(\frac{2m}{\Omega\gamma}\right).$
By substituting this equation and Eq. (118b) into Eq. (115), we arrive at Eq. (121), which is shown at the bottom of this page. Eq. (121a) holds because, as demonstrated in [48], the following inequality holds:
$E_{1}(x)>\frac{1}{2}e^{-x}\ln\left(1+\frac{2}{x}\right),\quad x>0.$ (122)
Thus, we have completed the proof for the upper bound of $s_{\infty}$ for $m\geq 1$.
@@ D-A3 $1=\frac{m}{\theta NT\log_{2}e}$
$=\frac{m}{\theta NT\log_{2}e}.$
D-B Lower bound of $s_{\infty}$ for $m\geq 1$
Since $\sqrt{1-x}\leq 1$ holds for $x\in(0,1)$, the lower bound of $s_{\infty}$ in the Nakagami-$m$ fading channel is derived in Eq. (123), shown at the bottom of this page. Eq. (123b) holds since $m\geq 1$ and $\frac{1}{\gamma}+x>x$ for $x>0$. Since $\frac{m}{\Omega}$ is a finite constant, we have
$\lim\limits_{\gamma\to+\infty}\frac{m}{\Omega\gamma}=0.$
Based on Eq. (123), we will divide the following discussion into two parts regrading the value of $\frac{m}{\theta NT\log_{2}e}$.
D-B1 $1<\frac{m}{\theta NT\log_{2}e}$
$\Gamma(s,x)<\frac{(x+v_{s})^{s}-x^{s}}{v_{s}s}e^{-x},$ (124)
where $v_{s}=\Gamma(s+1)^{\frac{1}{s-1}}>0$. For simplicity, we define
$c_{7}=\Gamma(m+1-\theta NT\log_{2}e)^{\frac{1}{m-1-\theta NT\log_{2}e}}.$
Based on Eq. (124), for $m-\theta NT\log_{2}e\in(0,1]\cup[2,\infty)$, we obtain Eq. (125), which is presented at the bottom of this page. Eq. (125a) holds because
$\lim_{\gamma\to+\infty}\frac{\ln\left[\left(1+\frac{c_{7}\Omega}{m}\gamma\right)^{m-\theta NT\log_{2}e}-1\right]}{(m-\theta NT\log_{2}e)\ln\left(1+\frac{c_{7}\Omega}{m}\gamma\right)}=1,$ (126)
where $c_{7}>0$ and $m-\theta NT\log_{2}e>0$. Eq. (125b) holds because $\lim\limits_{\gamma\to+\infty}\frac{\ln\left(1+\frac{c_{7}\Omega}{m}\gamma\right)}{\ln\gamma}=1$.
For $m-\theta NT\log_{2}e\in(1,2)$, we focus on the integral in Eq. (123a). When $m-\theta NT\log_{2}e\in(1,2)$, the condition $\theta NT\log_{2}e<m$ holds. Therefore, we find that the integral $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ is bounded by a finite constant according to Eq. (116). Thus, we obtain Eq. (127), which is presented at the bottom of this page.
D-B2 $1\geq\frac{m}{\theta NT\log_{2}e}$
For $m-\theta NT\log_{2}e<0$, we obtain Eq. (128), which is presented at the bottom of the last page. Eq. (128a) holds according to Eq. (120).
D-B3 $1=\frac{m}{\theta NT\log_{2}e}$
For $m-\theta NT\log_{2}e=0$, we have
$\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right)=E_{1}\left(\frac{m}{\Omega\gamma}\right).$
Based on Eq. (123), we obtain Eq. (129), which is presented at the bottom of this page. Eq. (129a) holds because according to [48], the following inequality holds:
@@ D-B3 $1=\frac{m}{\theta NT\log_{2}e}$
Therefore, according to the squeeze theorem, we have completed the proof for $m\geq 1$ based on the results shown in Eqs. (117), (119), (121), (125), (128), and (129).
@@
For $10.5\leq m<$ 1, there are some differences from the previous proofs for $m\geq 1$. The differences mainly lie in the proof for the upper bound when $1>\frac{m}{\theta NT\log_{2}e}$ and the lower bound.
For the upper bound when $1\leq\frac{m}{\theta NT\log_{2}e}$, the proof follows the same reasoning as presented in Eqs. (117) and (121). In contrast, for the upper bound with $1>\frac{m}{\theta NT\log_{2}e}$, Eq. (118) is changed into
@@
Then, by substituting Eq. (131) into Eq. (115), we derive Eq. (132), which is presented at the bottom of this page. Eq. (132a) holds because $m-\theta NT\log_{2}e<0$, which allows the application of Eq. (120).
@@
$=$ $\gamma^{\theta NT\log_{2}e-m}e^{\frac{m}{2\Omega\gamma}}K_{0}\left(\frac{m}{2\Omega\gamma}\right).$
@@
$=\frac{m}{\theta NT\log_{2}e}.$ (139b)
D-D Lower Bound of $s_{\infty}$ for $m\in[0.5,1)$
Next, we focus on the lower bound of $s_{\infty}$ for $10.5\leq m<$ 1. The essence of this part of the proof remains the scaling of the integral $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$. According to Eq. (123), the lower bound of $s_{\infty}$ is given in Eq. (133), located at the bottom of this page.
D-D1 $1<\frac{m}{\theta NT\log_{2}e}$
For $1<\frac{m}{\theta NT\log_{2}e}$, we transform Eq. (133) into Eq. (134), which is presented at the bottom of the next page. Eq. (134) holds because $-\theta NT\log_{2}e<0$ and $\frac{1}{\gamma}+x>x$. As mentioned previously, $\int_{0}^{\infty}z^{a}e^{-x}{\rm d}x<\infty$ holds when $a>-1$. Since in this case $1<\frac{m}{\theta NT\log_{2}e}$, we have $-\theta NT\log_{2}e+m-1>-1$. Thus, in this case the following integral is a finite constant:
$\int_{0}^{\infty}x^{-\theta NT\log_{2}e+m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$
Consequently, we obtain
$s_{\infty}$ $\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}-$ (135)
$\quad\quad\quad\quad\frac{1}{N}\lim_{\gamma\to+\infty}\frac{\beta}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}$
$=1.$
$=$ $\left(\frac{c_{9}}{c_{9}+1}\right)^{\theta T\beta+m-1}\left(\frac{m}{\Omega}\right)^{\theta N\log_{2}e-m}e^{\frac{m}{\Omega\gamma}}\Gamma\left(m-\theta NT\log_{2}e,\frac{m(1+c_{9})}{\Omega\gamma}\right).$
D-D2 $1\geq\frac{m}{\theta NT\log_{2}e}$
For $1\geq\frac{m}{\theta NT\log_{2}e}$, we obtain Eq. (136), which is presented at the bottom of this page. In Eq. (136), $K_{v}(z)$ denotes the modified Bessel function of the second kind [48]. Eq. (136a) holds because in this case, we have $m-\frac{1}{2}\geq 0$ and $\frac{x}{\frac{1}{\gamma}+x}\leq 1$. According to [48], for the modified Bessel function of the second kind $K_{v}(z)$ with $v=0$, we obtain
$\lim_{z\to 0}\frac{K_{0}(z)}{-\ln(z)}=c_{8},$ (137)
where $c_{8}>0$ is a constant.
Based on this conclusion, we transformed Eq. (136) into Eq. (138), which is presented at the bottom of this page. By substituting Eq. (138) into Eq. (133), we obtain Eq. (139), located at the bottom of this page. Eq. (139b) holds according to Eq. (137). By applying the squeeze theorem, we have completed the proof for $m\in[0.5,1)$. Thus, we have finished the proof.
@@ Appendix E Proof of Theorem 6
With the $\Psi(\gamma)$ satisfying Theorem 3, Eq. (113) still holds. For the AWGN channel, we can easily obtain
@@
For an FBL-SISO system in a Nakagami-$m$ fading channel, we will prove two inequalities shown in Eqs. (141) and (142), which are located at the bottom of this page. In Eq. (142), $c_{3}>0$ is an arbitrary constant. Since $c_{3}>0$ can be an arbitrarily small constant, we obtain Eq. (50) based on Eqs. (141) and (142).
The derivation of Eq. (141) is based on the derivation of the lower bounds of $s_{\infty}$ in the proof of Theorem 5. For $m\geq 1$, we can directly obtain Eq. (141) according to Eqs. (125a), (128a), and (129b). For $m\in[0.5,1)$, we can derive Eq. (141) according to Eqs. (135) and (139a). Thus, we omit the details here.
The derivation of Eq. (142) is based on the derivation of the upper bounds of $s_{\infty}$. Different from the proof of Theorem 5, we start the proof with the help of $\sqrt{1-x}\geq 1-\sqrt{x}$ with $x\in(0,1)$. We obtain Eq. (143), which is presented at the bottom of this page.
Since $\frac{x}{x+1}\leq\ln(1+x)$ holds for $x>0$, we have
$\frac{1}{1+\gamma x}\leq\ln\left(1+\frac{1}{\gamma x}\right).$
Therefore, we can transform Eq. (143) into Eq. (144), which is presented at the bottom of this page. The integral in Eq. (144) can be further manipulated to obtain Eq. (145), which is also presented at the bottom of this page.
@@
Note that, as $\gamma\to+\infty$, we have $\beta\to+\infty$. Therefore, $01-m-\theta T\beta<$ 0 holds. Based on this condition, we derive Eq. (146) at the bottom of this page, where $c_{9}$ is an arbitrarily positive constant. Based on this result, we obtain Eq. (147), which is also provided at the bottom of this page.
For $1>\frac{m}{\theta NT\log_{2}e}$, by combining Eq. (120) with Eq. (147), we obtain Eq. (148), which is at the bottom of this page. Note that, since $c_{9}$ is an arbitrarily positive constant, $\ln\left(1+\frac{1}{c_{9}}\right)$ is also an arbitrarily positive constant.
E-1 $1\leq\frac{m}{\theta NT\log_{2}e}\leq 1+\frac{1}{\theta NT\log_{2}e}$ and $\frac{m}{\theta NT\log_{2}e}\geq 1+\frac{2}{\theta NT\log_{2}e}$
The first part is $1\leq\frac{m}{\theta NT\log_{2}e}\leq 1+\frac{1}{\theta NT\log_{2}e}$ and $\frac{m}{\theta NT\log_{2}e}\geq 1+\frac{2}{\theta NT\log_{2}e}$, which implies $0\leq m-\theta NT\log_{2}e\leq 1$ and $m-\theta NT\log_{2}e\geq 2$. We will utilize the conclusion that the following inequality holds for $x>0$ with $s\leq 1$ and $s\geq 2$:
$\Gamma(s,x)\geq e^{-x}(1+x)^{s-1}.$
This can be easily demonstrated using Jensen’s inequality, as $(t+x)^{s-1}$ is a convex function of $t$ with a $s\leq 1$ and $s\geq 2$. As a result, we have
$\int_{x}^{\infty}y^{s-1}e^{-y}{\rm d}y$ $=e^{-x}\int_{0}^{\infty}(t+x)^{s-1}e^{-t}{\rm d}t$ (149)
$\geq e^{-x}\left(\int_{0}^{\infty}(t+x)e^{-t}{\rm d}t\right)^{s-1}$
$=e^{-x}(1+x)^{s-1}.$
Now we have proved that $\Gamma(s,x)\geq e^{-x}(1+x)^{s-1}$ holds for $s\leq 1$ and $s\geq 2$. Therefore, by substituting Eq. (149) into Eq. (147), we obtain Eq. (150), which appears at the bottom of this page.
E-2 $1+\frac{1}{\theta NT\log_{2}e}<\frac{m}{\theta NT\log_{2}e}<1+\frac{2}{\theta NT\log_{2}e}$
The second part considers $1+\frac{1}{\theta NT\log_{2}e}<\frac{m}{\theta NT\log_{2}e}<1+\frac{2}{\theta NT\log_{2}e}$, which implies $21<m-\theta NT\log_{2}e<$ 2. For $\Gamma(s,x)$ with $21<s<$ 2, according to [49], we have
$\Gamma(s,x)>\frac{(x+v_{s})^{s}-x^{s}}{v_{s}s}e^{-x},$ (151)
where $v_{s}$ is defined after Eq. (124). By substituting Eq. (151) into Eq. (147), we can derive Eq. (152) using similar procedures as shown in Eq. (125). Eq. (152) is presented at the bottom of this page. Now we have completed the proof of Eq. (142).
$\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}$ $\geq\lim_{\gamma\to+\infty}\frac{\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}}{\sqrt{N}\log_{2}\gamma}-\lim_{\gamma\to+\infty}\frac{1}{Q^{-1}(\epsilon(\gamma))\cdot\sqrt{N}\log_{2}\gamma}$ (160)
$=\frac{c_{4}}{\log_{2}e}\varpi.$
[…]